# Urban Scaling is hardwired in the human brain

### v1

@@ Abstract
The emerging field of the Science of Cities has unveiled previously undiscovered facets of urban life. Contrary to the expectation of chaotic behaviour influenced solely by cultural and geographic factors, cities globally exhibit universal power-law trends in urban scaling. Leveraging recent advances in mathematical descriptions of urban dynamics, this study investigates the interplay among fundamental allometry, fractal dimension, and the number of close contacts within cities. Through a dynamic exploration of these factors, a causal relationship is established, shedding light on the intricate dynamics that shape urban environments. Remarkably, our analysis suggests that the Big Five personality traits play a pivotal role in determining the fractal aspects of urban life. This research contributes to the understanding of urban scaling and proposes a novel connection between human personality traits and the structural patterns observed in cities, opening avenues for further interdisciplinary exploration.
@@
The fundamental allometric relationship, linking urban infrastructure area ($A$) to population size ($N$), is expressed as
@@
The fractal geometry of urban areas can be observed through cartographic analyses. The box-counting technique accurately determines the fractal dimension ($d_{f}$) of cities [6]. If $\ell$ represents the city’s typical linear length, its area is given by $A=\ell^{d_{f}}$. The fractal dimension typically falls within the range $1.91.4<d_{f}<$ 1.9, which is smaller than the expected integer dimension $d=2$. Key questions arise: What is the relationship between the scaling exponent and the fractal dimension? Why do humans exhibit fractal-like behaviour?
Recent developments have provided a mathematical framework for describing the dynamics of systems that move through fractal spaces [7]. In these systems, dynamics follow a nonlinear equation where the entropic index from Tsallis Statistics [8] plays a crucial role. When applied to model population density dynamics in cities, this framework yields the following relationship [9]:
$\beta=1-(q-1)=1-\frac{d-d_{f}}{2}\,.$ (2)
These equations establish connections between $\beta$, $q$, and $\delta d_{f}=d-d_{f}$, where $d_{f}$ is the fractal dimension of a fractal space embedded in a space with integer dimension $d$. The entropic index, $q$, is a parameter in the nonlinear equation describing anomalous diffusion in fractal spaces, and in the nonextensive statistics as a measure of entropy non-additiveness.
@@
A city fundamentally operates as a mechanism for generating social outputs, a concept central to various urban scaling models [2, 4]. Recent analyses reinforce this perspective by identifying links between urban scaling and the percolation fractal exponent [9]. In percolation systems, a critical density of connections between nodes determines when the system becomes fully interconnected, spanning from one end to the other. When applied to cities, this suggests that once a critical threshold of connectivity is achieved, individuals can travel easily across the city. Consequently, cities tend to evolve toward this critical percolation threshold, undergoing an organic development process to enhance connectivity among inhabitants. However, there is no indication that individuals consciously navigate fractal paths or adjust their movements to produce a power-law output. In this context, the scaling exponent emerges as a byproduct of urban organization, shaped by the population and infrastructure dynamics designed to maximize social output. This implies that urban output scaling is not the primary cause of fractal characteristics in city structure; rather, population dynamics and fractal geometry likely play more fundamental roles in driving these patterns.
Urban planning traditionally relies on a rational approach to designing infrastructure, typically based on Euclidean geometry. In this context, fractal aspects are not initially considered. Moreover, in many cases, the actual configuration of urban infrastructure results from centuries of continuous development and organic growth, where calculated planning has played a limited role. Consequently, fractal geometry likely arises more from social needs than from rational design decisions. This suggests that the dynamic flow, characterized by the variable $q$, determines the fractal geometry with dimension, $d_{f}$, clarifying the causality relation in Eq. 2. These insights reveal that human behaviour is resilient enough to transcend Euclidean geometric constraints, fostering the emergence of fractal geometry in the organic development of cities.
@@
Studies on information spreading and the epidemic evolution of diseases [10] highlight the significance of social contact in information diffusion and social output production. These models assume a fractal structure of social contacts, which leads to nonlinear information diffusion and results in informational distributions with a nonextensive nature. Empirical studies indicate that the dynamical parameter $q$ is determined by the average number of close contacts ($n_{c}$) through the relation $q-1=1/n_{c}$. Using this relation alongside Eqs 1 and 2, it follows that
@@
The above considerations underscore the crucial role of the number of contacts in social interaction. This number, linked to the hierarchical structure of social behaviour, regulates the flow of information within society. Accordingly, this study clarifies the relationships among social dynamics, fractal geometry, and the scaling of social outputs. It examines the causal relationships within urban processes, suggesting that individual dynamics—particularly the number of close contacts—are the primary drivers of fractal urban geometry. However, one question remains: why is such behaviour universal?
@@
The accepted consensus suggests that the crucial personality traits influencing social interactions amount to five [19]. While these traits do exert some influence on the number of friends an individual may have, the observed variations are relatively minor[20]. The average number of close friends aligns with findings from other independent studies. It is reasonable to posit that, given the necessity to identify a limited number of traits, the human brain specializes in learning a minimal set of personality models rather than generating numerous models. Consequently, for fostering a robust social life, it is preferable to intimately know a select few close contacts (such as family and friends) rather than maintaining superficial connections with a large number of acquaintances.
The findings support the social brain hypothesis, highlighting the deep links between human brain structure and social behavior, particularly the number of close relationships an individual can maintain. Studies of social interactions suggest a fractal organization in human social structures, with individuals forming layered networks based on social closeness. This hierarchical, fractal pattern in social organization constrains how people move within urban spaces: individuals preferentially visit locations where they can encounter familiar faces and often travel with acquaintances. Since commuting behavior plays a key role in urban planning decisions, the fractal nature of social relationships significantly shapes the design and layout of cities. Consequently, urban areas often exhibit fractal geometry, and social dynamics within these spaces follow a power-law distribution.
@@
In conclusion, this work reviews recent advances in understanding the dynamics of urban life. While the framework in Ref. [9] focuses on the physical and mathematical modeling of urban scaling, this study bridges these findings with the neurological and cognitive mechanisms that underpin human social behavior, providing a deeper understanding of the universal drivers of urban scaling. The anomalous diffusion of the population through fractal urban spaces establishes a connection between the fractal dimension and the scaling exponent. This anomalous diffusion is characterized by a non-linear equation associated with fractal geometries and non-extensive statistics. Notably, the entropic index, $q$, serves as a key parameter governing anomalous diffusion, highlighting the relevance of Tsallis statistics in bridging natural and social sciences [23].

### v2

@@ Abstract
The emerging field of the Science of Cities has revealed universal power-law trends in urban scaling, transcending cultural and geographic variations. This study investigates the interplay between fundamental allometry, fractal dimensions, and city social dynamics. By linking these factors to the hierarchical social structures defined by Dunbar’s numbers, a causal relationship is established that illuminates the drivers of urban scaling. Notably, this research highlights the pivotal role of the Big Five personality traits in shaping the fractal nature of urban environments. These findings bridge the gap between human cognitive traits and urban structural patterns, offering a novel perspective on the universality of urban scaling and opening new opportunities for interdisciplinary exploration.
@@
The fundamental allometric relationship, linking urban infrastructure area ($A$) to population size ($N$), is expressed as [^1] The allometric relation connects the urban area to population size and is inspired by the metabolic scaling relationships observed in mammals.
@@
The fractal geometry [^2] Fractals are mathematical or natural structures characterized by scaling invariance and fine internal details. In geometry, these properties often result in spaces with fractional dimensions. A key feature of fractals is self-similarity, meaning that a portion of the structure resembles the entire structure when magnified. of urban areas can be observed through cartographic analyses. The box-counting technique accurately determines the fractal dimension ($d_{f}$) of cities [6]. If $\ell$ represents the city’s typical linear length, its area is given by $A=\ell^{d_{f}}$. The fractal dimension typically falls within the range $1.91.4<d_{f}<$ 1.9, which is smaller than the expected integer dimension $d=2$. Key questions arise: What is the relationship between the scaling exponent and the fractal dimension? Why do humans exhibit fractal-like behaviour?
Recent developments have provided a mathematical framework for describing the dynamics of systems that move through fractal spaces [7]. In such systems, dynamics follow a nonlinear equation where the entropic index from Tsallis Statistics [8] plays a crucial role [^3] Tsallis Statistics is a generalization of Boltzmann Statistics, proposed in 1988 to address the growing number of systems that standard statistical mechanics could not appropriately describe. In particular, this generalized framework is especially effective in modelling complex systems.. When applied to model population density dynamics in cities, this framework yields the following relationship [9]:
$\beta=1-\frac{d}{2}(q-1)=1-\frac{d-d_{f}}{2}\,.$ (2)
[…]
@@
A city fundamentally operates as a mechanism for generating social outputs, a concept central to various urban scaling models [2, 4]. Recent analyses reinforce this perspective by identifying links between urban scaling and the percolation fractal exponent [9]. In percolation systems, a critical density of connections between nodes determines when the system becomes fully interconnected, spanning from one end to the other [^5] In the allometric models, a city’s infrastructure, similar to the circulatory system of mammals, is built to ensure that all individuals have access to any place within the city, forming a percolative system.. When applied to cities, this suggests that once a critical threshold of connectivity is achieved, individuals can travel easily across the city. Consequently, cities tend to evolve toward this critical percolation threshold, undergoing an organic development process to enhance connectivity among inhabitants. However, there is no indication that individuals consciously navigate fractal paths or adjust their movements to produce a power-law output. In this context, the scaling exponent emerges as a byproduct of urban organization, shaped by the population and infrastructure dynamics designed to maximize social output. This implies that urban output scaling is not the primary cause of fractal characteristics in city structure; rather, population dynamics and fractal geometry likely play more fundamental roles in driving these patterns.
Urban planning traditionally relies on a rational approach to designing infrastructure typically based on Euclidean geometry. In this context, fractal aspects are not initially considered. Moreover, in many cases, the actual configuration of urban infrastructure results from centuries of continuous development and organic growth, where calculated planning has played a limited role. Consequently, fractal geometry likely arises more from social needs than from rational design decisions. This suggests that the dynamic flow, characterized by the variable $q$, determines the fractal geometry with dimension, $d_{f}$, clarifying the causality relation in Eq. 2. These insights reveal that human behaviour is resilient enough to transcend Euclidean geometric constraints, fostering the emergence of fractal geometry in the organic development of cities.
@@
Studies on information spreading and the epidemic evolution of diseases [10, 11] highlight the significance of social contact in information diffusion and social output production. These models assume a fractal structure of social contacts, which leads to nonlinear information diffusion and results in informational distributions with a nonextensive nature. The models indicate that the dynamical parameter $q$ is determined by the average number of close contacts ($n_{c}$) through the relation $q-1=1/n_{c}$ [^6] This relation was validated by an analysis of COVID-19 spread presented in Ref. [11]. Using the relation between $q$ and $n_{c}$ along with Eqs 1 and 2, it follows that
@@
The above considerations emphasize the critical role of the number of contacts in social interactions. This number, linked to the hierarchical structure of social behaviour, governs the flow of information within society by determining the nonlinearity of the social dynamics regulated by the parameter $q$ in the equation describing the dynamic process. Consequently, this study clarifies the relationships among key social aspects: the number of close contacts, social dynamics within a hierarchical structure, fractal geometry, and the scaling of social outputs. It explores the causal relationships underlying urban processes, suggesting that individual behaviour —particularly the number of close contacts — serves as the primary driver of fractal urban geometry. Yet, one fundamental question remains: why is such behaviour universal?
@@
The accepted consensus suggests that the crucial personality traits influencing social interactions are five in number [20]. These fundamental traits, collectively referred to as the Big Five, are believed to encompass the essential social characteristics that individuals must evaluate to maintain a healthy social life [^7] The Big Five personality traits are: Openness, Conscientiousness, Extraversion, Agreeableness and Neuroticism.. While these traits do exert some influence on the number of friends an individual may have, the observed variations are relatively minor[21]. The average number of close friends aligns with findings from other independent studies. It is reasonable to posit that, given the necessity to identify a limited number of traits, the human brain specializes in learning a minimal set of personality models rather than generating numerous models. Consequently, for fostering a robust social life, it is preferable to intimately know a select few close contacts (such as family and friends) rather than maintaining superficial connections with a large number of acquaintances.
The findings support the social brain hypothesis, highlighting the deep links between human brain structure and social behaviour, particularly the number of close relationships an individual can maintain [^8] The hierarchical social structure was proposed as an extension of Dunbar’s number, which represents the limit on the number of friends a person can maintain—approximately 150. Further analysis by Dunbar and colleagues revealed that the group of 150 friends exhibits non-uniformities in the types of social interactions, leading to a hierarchical structure that is approximately scale-free, similar to a fractal structure.. Studies of social interactions suggest a fractal organization in human social structures, with individuals forming layered networks based on social closeness. This hierarchical, fractal pattern in social organization constrains how people move within urban spaces: individuals preferentially visit locations where they can encounter familiar faces and often travel with acquaintances. Since commuting behaviour plays a key role in urban planning decisions, the fractal nature of social relationships significantly shapes the design and layout of cities. Consequently, urban areas often exhibit fractal geometry, and social dynamics within these spaces follow a power-law distribution.
@@
In conclusion, this work reviews recent advances in understanding the dynamics of urban life. While the framework in Ref. [9] focuses on the physical and mathematical modelling of urban scaling, this study bridges these findings with the neurological and cognitive mechanisms that underpin human social behaviour, providing a deeper understanding of the universal drivers of urban scaling. The anomalous diffusion of the population through fractal urban spaces establishes a connection between the fractal dimension and the scaling exponent. This anomalous diffusion is characterized by a non-linear equation associated with fractal geometries and non-extensive statistics. Notably, the entropic index, $q$, serves as a key parameter governing anomalous diffusion, highlighting the relevance of Tsallis statistics in bridging natural and social sciences [24].